# Scaling Up Membership Inference: When and How Attacks Succeed on Large Language Models

### v1

@@ 1 Introduction
The application of MIAs on LLMs has faced discouraging results so far. Initial claims of success (Shi et al., 2024; Meeus et al., 2024a; Carlini et al., 2021; Mattern et al., 2023) were later disproven by Duan et al. (2024); Das et al. (2024); Maini et al. (2024); Meeus et al. (2024b). Before their work, it was common practice to select true non-members from documents created after the LLM’s cut-off date. They showed that this approach allowed MIA methods to exploit temporal cues, rather than identifying membership through the model’s inherent response characteristics. Duan et al. (2024); Das et al. (2024); Maini et al. (2024) introduced a new evaluation method based on an independent, identically distributed (IID) split between true members and non-members. Since adopting this method, no further MIA studies on LLMs have shown performance significantly better than random. Reported membership detection has remained below 60% AUROC, close to the random-chance level of 50% AUROC (Duan et al., 2024; Das et al., 2024; Maini et al., 2024; Xie et al., 2024; Zhang et al., 2024b).
@@ 3 MIA Evaluation
Sentence-level MIA. We define a sentence as a natural sequence of words ending in a full stop. The average sentence in the Pile contains 43 tokens. This granularity is important because it is used to detect whether specific data points in benchmarks (e.g., questions in question-answering tasks) are contaminated, ensuring fair model evaluation. Due to its short nature, it can be extremely challenging to perform MIA successfully. For instance, Duan et al. (2024) shows large overlaps between member and non-member sentences, which blurs the decision boundary. Additionally, sentence-level MIA can probe privacy leakage inferring membership of personally identifiable information (Kim et al., 2023).
Paragraph-level MIA. We define a paragraph as a sequence of tokens that fits within the context window of a large language model (LLM). Thus, the length of a "paragraph" depends on the specific LLM in use. In our experiments, we use paragraphs of up to 512, 1024, and 2048 tokens, aligning with the context window sizes of the target models. In practice, paragraph-level copyrighted materials are particularly relevant to user-generated content on social media platforms and online forums, where short-form content is most common.
Document-level MIA. We define a document in the conventional sense, such as a single arXiv paper. In The Pile dataset, the average document contains 14,222 tokens. Documents typically consist of multiple paragraphs, often exceeding the context window of many LLMs, such as Pythia, which has a 2048-token limit. MIA approaches must handle these long documents by splitting the token sequence into smaller chunks, or “paragraphs” (as defined earlier), that fit within the model’s context length and then aggregating the model’s responses across these chunks. Performing MIA at the document level is essential for addressing copyright and data ownership concerns, particularly for copyrighted materials like novels, news articles, research papers, and book chapters.
Collection-level MIA. We define a collection as a set of documents. The attacker can choose the collection’s size. For instance, with 100 documents, a collection contains approximately 1.4 million tokens. Collection-level MIA is crucial, as one may question whether a collection of articles, such as those provided by an internet service provider, has been crawled and used in LLM training. Aggregating individual signals can amplify the detection of the collection’s usage. This is also relevant when examining contaminated benchmarks, as entire datasets may be unintentionally included in the training, leading to an overestimation of model performance.
@@
Pretraining MIA requires the attacker to identify whether a given data instance was part of the pretraining corpus.
Continual-learning MIA addresses the challenge of keeping models up-to-date, as frequently retraining them from scratch is often infeasible due to high costs. Instead, continual learning offers a solution by incrementally training the model’s checkpoint on new data. This approach is crucial for ensuring that LLMs remain current with evolving information (Wu et al., 2024).
Fine-tuning MIA seeks to determine whether specific samples were used during fine-tuning for particular end tasks or domains, such as training data for conversational bots, question-answering systems, and similar applications. Fine-tuning methods are becoming increasingly popular due to their lower costs and flexibility, making them an ideal solution for tasks with large amounts of relevant data (Meta, 2024).
@@ 4 Method
After aggregation, we obtain a set of $K$ MIA scores: $\{f(t_{p,1}),\cdots,f(t_{p,K})\}$. These scores are treated as an unordered set, as the criteria for determining membership becomes largely independent of the ordering of documents within a dataset or paragraphs within a document. Following the DI approach, we perform statistical testing by comparing the set of $K$ MIA scores against a baseline set of $K$ scores from known non-members. Specifically, we use the Student’s t-test for collection-level MIA since we aggregate more than 30 samples and the Mann–Whitney U-test for document-level MIA. since we aggregate less than 30 samples from a non-normal distribution.
@@
Here, $\mu$ and $s$ represent the mean and sample standard deviation of the $K$ MIA scores, with the subscript n-m indicating the scores from the non-members. The term $\text{rank}(f(t_{p,k}))$ refers to the rank of the MIA feature for the query sequence $t$ in the combined set of $2K$ MIA features, which includes both the query sequence and the known non-members.
We conduct our main experiments on Pythia 2.8B and 6.9B (Biderman et al., 2023) with data from The Pile dataset (Gao et al., 2020). See Appendix A for details on the experimental setup.
@@ 5 Experiments
Figure 4: Failutre modes for aggregation. Aggregation may not work when the base performance is too low (left and right plots), or when the amount of information to aggregate is too short (centre plot).
@@ 5.1 Aggregating Text Subunits is Effective
The left plot focuses on collection MIA for arXiv. It reveals a stark contrast between small and large collections. While collections with only dozens of documents yield AUROC scores barely above random chance, those with 500 or more documents achieve significantly higher AUROC. Notably, using the 6.9B model and despite employing only a few MIA methods in our ensemble, we attain a remarkable collection MIA AUROC of 0.9 for arXiv.
The right plot shows the performance of document MIA on arXiv, where we observe our highest results. This exceptional performance can be attributed to two factors: i) the considerable length of the documents (averaging around 15K tokens), allowing for the aggregation of numerous MIA scores, and ii) a paragraph MIA AUROC exceeding 0.53. The combination of these factors yields an impressive document MIA AUROC of 0.75. To the best of our knowledge, this marks the first successful application of MIA to entire documents.
@@ Appendix A Experimental Settings
We use the MIA implementations from Shi et al. (2024). We run all our experiments on 5 random seeds. We report the average AUROC with its standard deviation.
@@ Appendix B MIA Benchmmark
Dataset Model Collection MIA Document MIA Paragraph MIA Sentence MIA
ArXiv 2.8B 0.718 ± 0.122 0.523 ± 0.01 0.509 ± 0.006 0.499 ± 0.003
6.9B 0.943 ± 0.025 0.697 ± 0.06 0.557 ± 0.008 0.501 ± 0.003
FreeLaw 6.9B 0.551 ± 0.126 0.538 ± 0.024 0.526 ± 0.011 0.505 ± 0.001
2.8B 0.511 ± 0.106 0.51 ± 0.017 0.509 ± 0.014 0.504 ± 0.002
Github 6.9B 0.497 ± 0.088 0.491 ± 0.005 0.494 ± 0.006 0.865 ± 0.004
2.8B 0.479 ± 0.069 0.498 ± 0.01 0.494 ± 0.009 0.496 ± 0.006
HackerNews 6.9B 0.709 ± 0.34 0.513 ± 0.04 0.514 ± 0.018 0.5 ± 0.003
2.8B 0.686 ± 0.267 0.488 ± 0.02 0.511 ± 0.025 0.498 ± 0.002
OpenWebText2 2.8B 0.544 ± 0.097 0.505 ± 0.009 0.502 ± 0.008 0.497 ± 0.003
6.9B 0.615 ± 0.089 0.519 ± 0.008 0.513 ± 0.006 0.499 ± 0.005
Pile-CC 2.8B 0.52 ± 0.095 0.489 ± 0.011 0.481 ± 0.005 0.495 ± 0.004
6.9B 0.55 ± 0.132 0.513 ± 0.015 0.5 ± 0.013 0.931 ± 0.005
USPTO 2.8B 0.512 ± 0.112 0.495 ± 0.032 0.49 ± 0.014 0.5 ± 0.001
6.9B 0.609 ± 0.151 0.534 ± 0.007 0.52 ± 0.005 0.497 ± 0.002
Wikipedia 6.9B 0.85 ± 0.144 0.56 ± 0.011 0.553 ± 0.01 0.507 ± 0.004
2.8B 0.665 ± 0.169 0.531 ± 0.019 0.534 ± 0.015 0.503 ± 0.006

### v2

@@ 1 Introduction
The application of MIAs on pretrained LLMs has faced discouraging results so far. Despite the initial optimism from the works of (Shi et al., 2024; Meeus et al., 2024a; Carlini et al., 2021; Mattern et al., 2023), Duan et al. (2024); Das et al. (2024); Maini et al. (2024); Meeus et al. (2024b) show later that current methods only achieve near random-guessing performance. Before their work, it was common practice to select true non-members from documents created after the LLM’s cut-off date. They showed that this approach allowed MIA methods to exploit temporal cues, rather than identifying membership through the model’s inherent response characteristics. Duan et al. (2024); Das et al. (2024); Maini et al. (2024) introduced a new evaluation method based on an independent, identically distributed (IID) split between true members and non-members. Since adopting this method, no further MIA studies on LLMs have shown performance significantly better than random. Reported membership detection has remained below 60% AUROC, close to the random-chance level of 50% AUROC (Duan et al., 2024; Das et al., 2024; Maini et al., 2024; Xie et al., 2024; Zhang et al., 2024b).
@@ 3 MIA Evaluation
1) Sentence-level MIA. We define a sentence as a natural sequence of words ending in a full stop. The average sentence in the Pile contains 43 tokens. This granularity is important because it is used to detect whether specific data points in benchmarks (e.g., questions in question-answering tasks) are contaminated, ensuring fair model evaluation. Due to its short nature, it can be extremely challenging to perform MIA successfully. For instance, Duan et al. (2024) shows large overlaps between member and non-member sentences, which blurs the decision boundary. Additionally, sentence-level MIA can probe privacy leakage inferring membership of personally identifiable information (Kim et al., 2023).
2) Paragraph-level MIA. We define a paragraph as a sequence of tokens that fits within the context window of a large language model (LLM). Thus, the length of a "paragraph" depends on the specific LLM in use. In our experiments, we use paragraphs of up to 512, 1024, and 2048 tokens, aligning with the context window sizes of the target models. In practice, paragraph-level copyrighted materials are particularly relevant to user-generated content on social media platforms and online forums, where short-form content is most common.
3) Document-level MIA. We define a document in the conventional sense, such as a single arXiv paper. In The Pile dataset, the average document contains 14,222 tokens. Documents typically consist of multiple paragraphs, often exceeding the context window of many LLMs, such as Pythia, which has a 2048-token limit. MIA approaches must handle these long documents by splitting the token sequence into smaller chunks, or “paragraphs” (as defined earlier), that fit within the model’s context length and then aggregating the model’s responses across these chunks. Performing MIA at the document level is essential for addressing copyright and data ownership concerns, particularly for copyrighted materials like novels, news articles, research papers, and book chapters.
4) Collection-level MIA. We define a collection as a set of documents. The attacker can choose the collection’s size. For instance, with 100 documents, a collection contains approximately 1.4 million tokens. Collection-level MIA is crucial, as one may question whether a collection of articles, such as those provided by an internet service provider, has been crawled and used in LLM training. Aggregating individual signals can amplify the detection of the collection’s usage. This is also relevant when examining contaminated benchmarks, as entire datasets may be unintentionally included in the training, leading to an overestimation of model performance.
@@
1) Pretraining MIA requires the attacker to identify whether a given data instance was part of the pretraining corpus.
2) Continual-learning MIA addresses the challenge of keeping models up-to-date, as frequently retraining them from scratch is often infeasible due to high costs. Instead, continual learning offers a solution by incrementally training the model’s checkpoint on new data. This approach is crucial for ensuring that LLMs remain current with evolving information (Wu et al., 2024).
3) Fine-tuning MIA seeks to determine whether specific samples were used during fine-tuning for particular end tasks or domains, such as training data for conversational bots, question-answering systems, and similar applications. Fine-tuning methods are becoming increasingly popular due to their lower costs and flexibility, making them an ideal solution for tasks with large amounts of relevant data (Meta, 2024).
@@ 4 Method
During the second stage, we obtain a set of $K$ MIA scores: $\{f(t_{p,1}),\cdots,f(t_{p,K})\}$. These scores are treated as an unordered set, as the criteria for determining membership becomes largely independent of the ordering of documents within a dataset or paragraphs within a document. Following the DI approach, we perform statistical testing by comparing the set of $K$ MIA scores against a baseline set of $K$ scores from known non-members. Specifically, we use the Student’s t-test for collection-level MIA since we aggregate more than 30 samples and the Mann–Whitney U-test for document-level MIA since we aggregate less than 30 samples from a non-normal distribution.
@@
Here, $\mu$ and $s$ represent the mean and sample standard deviation of the $K$ MIA scores, with the subscript n-m indicating the scores from the non-members. The term $\text{rank}(f(t_{p,k}))$ refers to the rank (i.e., position in the list) of the MIA feature for the query sequence $t$ in the combined set of $2K$ MIA features, which includes both the query sequence and the known non-members.
We conduct all our experiments on Pythia 2.8B and 6.9B (Biderman et al., 2023) with data from The Pile dataset (Gao et al., 2020). We also provide the results of our main experiments on GPT Neo 2.7B in Appendix B. See Appendix A for details on the experimental setup.
@@ 5 Experiments
Figure 4: Failute modes for aggregation. Aggregation may not work when the base performance is too low (left and right plots), or when the amount of information to aggregate is too short (centre plot).
@@ 5.1 Aggregating Text Subunits is Effective
The upper plot focuses on collection MIA for arXiv. It reveals a stark contrast between small and large collections. While collections with only dozens of documents yield AUROC scores barely above random chance, those with 500 or more documents achieve significantly higher AUROC. Notably, using the 6.9B model and despite employing only a few MIA methods in our ensemble, we attain a remarkable collection MIA AUROC of 0.9 for arXiv.
The bottom plot shows the performance of document MIA on arXiv, where we observe our highest results. This exceptional performance can be attributed to two factors: i) the considerable length of the documents (averaging around 15K tokens), allowing for the aggregation of numerous MIA scores, and ii) a paragraph MIA AUROC exceeding 0.53. The combination of these factors yields an impressive document MIA AUROC of 0.75. To the best of our knowledge, this marks the first successful application of MIA to entire documents.
Table 7 in Appendix B includes all the results of the membership inference attacks (MIA) across all data levels for GPT Neo 2.7B, Pythia 2.8B, and Pythia 6.9B, showing that the trends holds for the three models.
@@ Appendix A Experimental Settings
We use the MIA implementations from Shi et al. (2024). We run all our experiments on five random seeds. We report the average AUROC with its standard deviation. To train the linear mapping that aggregates multiple MIA signals in the first stage of our method (Section 4), we use 1k members and 1k non-members from the known partition.
Sentence-Level MIA.
Since our base unit for aggregations is the sentence, sentence-level MIA does not use the second stage of our method (i.e., the aggregation part in Section 4), and instead only uses the first stage. Hence, it consists of learning a linear classifier that aggregates all the MIA methods and assigns a score to the input sentence.
@@ Appendix B MIA Benchmmark
Dataset Model Collection Document Paragraph Sentence
ArXiv GPT-Neo 2.7B 0.608 ± 0. 163 0.522 ± 0.011 0.504 ± 0.006 0.497 ± 0.002
Pythia 2.8B 0.718 ± 0.122 0.523 ± 0.01 0.509 ± 0.006 0.499 ± 0.003
Pythia 6.9B 0.943 ± 0.025 0.697 ± 0.06 0.557 ± 0.008 0.501 ± 0.003
FreeLaw GPT-Neo 2.7B 0.483 ± 0.148 0.522 ± 0.011 0.498 ± 0.008 0.506 ± 0.002
Pythia 2.8B 0.511 ± 0.106 0.51 ± 0.017 0.509 ± 0.014 0.504 ± 0.002
Pythia 6.9B 0.551 ± 0.126 0.538 ± 0.024 0.526 ± 0.011 0.505 ± 0.001
Github GPT-Neo 2.7B 0.496 ± 0.058 0.500 ± 0.008 0.498 ± 0.003 0.497 ± 0.004
Pythia 2.8B 0.479 ± 0.069 0.498 ± 0.01 0.494 ± 0.009 0.496 ± 0.006
Pythia 6.9B 0.497 ± 0.088 0.491 ± 0.005 0.494 ± 0.006 0.865 ± 0.004
HackerNews GPT-Neo 2.7B 0.774 ± 0.077 0.511 ± 0.027 0.509 ± 0.007 0.501 ± 0.002
Pythia 2.8B 0.686 ± 0.267 0.488 ± 0.02 0.511 ± 0.025 0.498 ± 0.002
Pythia 6.9B 0.709 ± 0.34 0.513 ± 0.04 0.514 ± 0.018 0.5 ± 0.003
OpenWebText2 GPT-Neo 2.7B 0.518 ± 0.078 0.498 ± 0.015 0.5010 ± 0.003 0.497 ± 0.006
Pythia 2.8B 0.544 ± 0.097 0.505 ± 0.009 0.502 ± 0.008 0.497 ± 0.003
Pythia 6.9B 0.615 ± 0.089 0.519 ± 0.008 0.513 ± 0.006 0.499 ± 0.005
Pile-CC GPT-Neo 2.7B 0.516±0.089 0.497 ± 0.007 0.501 ±0.006 0.499 ± 0.007
Pythia 2.8B 0.52 ± 0.095 0.489 ± 0.011 0.481 ± 0.005 0.495 ± 0.004
Pythia 6.9B 0.55 ± 0.132 0.513 ± 0.015 0.5 ± 0.013 0.931 ± 0.005
USPTO GPT-Neo 2.7B 0.501 ±0.066 0.482±0.007 0.499±0.004 0.496±0.002
Pythia 2.8B 0.512 ± 0.112 0.495 ± 0.032 0.49 ± 0.014 0.5 ± 0.001
Pythia 6.9B 0.609 ± 0.151 0.534 ± 0.007 0.52 ± 0.005 0.497 ± 0.002
Wikipedia GPT-Neo 2.7B 0.648 ± 0.103 0.517 ± 0.017 0.508 ± 0.006 0.497 ± 0.005
Pythia 2.8B 0.665 ± 0.169 0.531 ± 0.019 0.534 ± 0.015 0.503 ± 0.006
Pythia 6.9B 0.85 ± 0.144 0.56 ± 0.011 0.553 ± 0.01 0.507 ± 0.004